OpenStudy (anonymous):

Is this very easy? Given that p, q are prime numbers such that \(p^2+q = 37q^2+p\). find \( p−q\) ?

OpenStudy (anonymous):

This problem is added only on the request of Mr.Math.

OpenStudy (anonymous):

p=43 and q=7 p-q=43-7=36

OpenStudy (mr.math):

The difference between two primes (none of which is 2) is an even number. Someone can easily see that 2 can't be a solution for either p nor q. So let \(p-q=2k, k\in \mathbb Z\), then: \[p^2-37q^2=2k \implies 2k(p+q)-36q^2=2k\] \[ \implies 2k=\frac{36q^2}{p+q-1} \implies k=\frac{18q^2}{p+q-1}.\] Sice k is an integer then \((p+q-1)|18q^2 \implies (p+q-1)|18 \text{ or }(p+q-1)|q^2.\) For (p+q-1)|18, there are few cases here in which none is correct. If (p+q-1)|q^2, then p+q-1 is equal 1, q or q^2. It's very obvious that the only possible case is that \(p+q-1=q^2\). Thus \(k=\frac{18q^2}{q^2}=18 \implies 2k=p-q=36\).

OpenStudy (anonymous):

If you are assuming \( p-q=2k, k\in \mathbb Z \), how can you say that \( p^2-37q^2=2k \) same 2k ?

OpenStudy (mr.math):

\[p^2+q=37q^2+p \implies p^2-37q^2=p-q=\cdots\]

OpenStudy (anonymous):

Yes good, you should include every step, you know that I am a FOOL :P

OpenStudy (anonymous):

Not bad Mr.Math ;)

OpenStudy (mr.math):

Thanks Mr.Fool ;D

OpenStudy (asnaseer):

Neat solution Mr.Math

OpenStudy (mr.math):

Thanks asnaseer! :-)

OpenStudy (mr.math):

I guess I'm learning from you :)

OpenStudy (asnaseer):

:-) - I doubt that Sir! You are far cleverer than you think.